# Dependent Dirichlet processes via thinning

### v1

@@ B.1 Proof of Proposition 4
We want to prove the full support of the thinned-DDP, where the full support is defined in the sense of Barrientos et al. (2012). In Theorem 1, they demonstrate that a sufficient condition to have full weak support is to ensure that the process assigns positive probability to product space of particular simplices. In the following, we show that such a condition holds for the thinned-DDP.
We study the setting where we only have two groups, but the proof can be easily generalized to any $G$. The sufficient condition of Barrientos et al. (2012) requires that
@@
where $A_{1},\dots,A_{k}$ is a measurable partition of $\Theta$, and the $q_{j,g}$ are constants such that $q_{j,g}\geq 0$, $\sum_{j=1}^{k}q_{j,g}=1$ for $g=1,2$.
Without loss of generality, we can assume that $P_{0}(A_{j})>0$ for $j=1,\dots,k$, since if $P_{0}(A_{j})=0$ then $p_{g}(A_{j})=0$ a.s. and $q_{j,g}=0$. We need to find a subset of $\Omega_{0}$ with positive probability. Assume that the sequences $\{\ell_{j,g}\}_{j\geq 1}$, $\{\theta_{j}\}_{j\geq 1}$ and $\{v_{j}\}_{j\geq 1}$ are independent. Consider the set
[…]
@@
while $\ell_{j,1}\in\{0,1\}$ for $j>k$, and $\ell_{j,2}\in\{0,1\}$ for $j>2k$, without restrictions. We assume that $\mathbb{P}(\Omega_{1})>0$ since this is true under all the considered prior distributions on the thinning variables.
Consider now
@@
$B_{3}$ is an open subset of $S^{k}$, at least for sufficiently small $\epsilon$. On $\Omega_{1}$ we have $\bm{w}_{k}=f(v_{1},\dots,v_{k})$, with $w_{1,1}=v_{1}$, $w_{j,1}=v_{j}\prod_{h<j}(1-v_{h})$: this mapping is one-to-one and continuous from $(0,1)^{k}$ to $S^{k}$. It follows that $f^{-1}(B_{3})=C_{3}$ is an open subset of $(0,1)^{k}$. Therefore,
@@ B.2 Proof of Proposition 5
×$\quad\times\sum_{j_{2}=j_{1}+1}q\prod_{i_{2}=j_{1}+1}^{j_{2}-1}\frac{\alpha+1-q}{\alpha+1}$
@@ B.3 Proof of Proposition 6
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{u_{1},u_{2}}\left[\mathrm{Corr}(p_{1}(A),p_{2}(A)\mid u_{1},u_{2})\sqrt{\mathrm{Pr}(y_{i,1}=y_{i^{\prime},1}\mid u_{1},u_{2})\mathrm{Pr}(y_{k,2}=y_{k^{\prime},2}\mid u_{1},u_{2})}\right]$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}_{u_{1},u_{2}}\left[\mathrm{Corr}(p_{1}(A),p_{2}(A)\mid u_{1},u_{2})\sqrt{\mathrm{Pr}(y_{i,1}=y_{i^{\prime},1})\mathrm{Pr}(y_{k,2}=y_{k^{\prime},2})}\right]$
@@ Appendix D Correlations with parent process
[…]
@@ Appendix F Additional details on the simulation studies
Estimation of the thinned-DDP was performed using the Gibbs sampler outlined in Section 4; for the complete-pooling and no-pooling models, we used a standard blocked Gibbs sampler for DP mixtures. The truncation parameter was set to 100 for all models. Posterior inference for the CAM and the GM-DDP was performed using the R packages SANple (D’Angelo & Denti 2023) and BNPmix (Corradin et al. 2021), respectively. We ran all algorithms for 3,000 iterations, discarding the first 2,000 as burn-in. Figure F1 shows the distribution of the computation time (in logarithmic scale) for the thinned-DDP, the CAM, and the GM-DDP.
@@ F.1 Additional figures
Figure F3 displays a summary of the distribution of the average length of the pointwise 95% highest posterior density (HPD) credible intervals for the thinned-DDP, the complete pooling, and no pooling mixtures. Figure F3 displays the same quantity for the thinned-DDP, the GM-DDP, and the CAM.

### v2

@@ B.1 Proof of Proposition 4
We want to prove that the weak support of the thinned-DDP is $\mathcal{P}(\Theta)^{G}$. It is simple to check that $\mathcal{P}(\Theta)^{G}$ is a closed set with probability measure one (see the proof of Theorem 1 in Barrientos et al.2012). It is left to show that $\mathcal{P}(\Theta)^{G}$ is the smallest closed set with probability one. In the proof of Theorem 1, Barrientos et al. (2012) demonstrate that a general sufficient condition for this is to ensure that the process assigns positive probability to certain basic open sets of $\mathcal{P}(\Theta)^{G}$, which can be written as a product space of particular subsets of simplices. In the following, we show that such a condition holds for the thinned-DDP.
We study the setting where we only have two groups, but the proof can be easily generalized to any $G$. Specifically, the sufficient condition of Barrientos et al. (2012) (see their formula (3)) can be reformulated as
@@
where $A_{1},\dots,A_{k}$ is a $P_{0}$-continuosus measurable partition of $\Theta$, and the $q_{j,g}$ are constants such that $q_{j,g}\geq 0$, $\sum_{j=1}^{k}q_{j,g}=1$ for $g=1,2$. Furthermore, we only have to consider sets $A_{j}$’s such that $P_{0}(A_{j})>0$ for $j=1,\dots,k$, since if $P_{0}(A_{j})=0$ then $p_{g}(A_{j})=0$ a.s. and $q_{j,g}=0$. We need to find a subset $\Omega_{0}$ with positive probability. Consider the set
$\Omega_{1}=\{$ $\ell_{j,1}=1\>\text{for $j=1,\dots,k$},$ (2)
@@
while $\ell_{j,1}\in\{0,1\}$ for $j>k$, and $\ell_{j,2}\in\{0,1\}$ for $j>2k$, without restrictions.
By assumption $\mathbb{P}(\Omega_{1})>0$. Consider now
@@
$B_{3}$ is an open subset of $S^{k}$. On $\Omega_{1}$ we have $\bm{w}_{k}=f(v_{1},\dots,v_{k})$, with $w_{1,1}=v_{1}$, $w_{j,1}=v_{j}\prod_{h<j}(1-v_{h})$: this mapping is one-to-one and continuous from $(0,1)^{k}$ to $S^{k}$. It follows that $f^{-1}(B_{3})=C_{3}$ is an open subset of $(0,1)^{k}$. Therefore,
@@ B.2 Proof of Proposition 5
$\quad\times\sum_{j_{2}=j_{1}+1}q\prod_{i_{2}=j_{1}+1}^{j_{2}-1}\frac{\alpha+1-q}{\alpha+1}$
@@ B.3 Proof of Proposition 6
$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{u_{1},u_{2}}\left[\mathrm{Corr}(p_{1}(A),p_{2}(A)\mid u_{1},u_{2})\sqrt{\mathrm{Pr}(y_{i,1}=y_{i^{\prime},1}\mid u_{1},u_{2})\mathrm{Pr}(y_{k,2}=y_{k^{\prime},2}\mid u_{1},u_{2})}\right]$
$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}_{u_{1},u_{2}}\left[\mathrm{Corr}(p_{1}(A),p_{2}(A)\mid u_{1},u_{2})\sqrt{\mathrm{Pr}(y_{i,1}=y_{i^{\prime},1})\mathrm{Pr}(y_{k,2}=y_{k^{\prime},2})}\right]$
@@ Appendix D Correlations with parent process
We consider a vector $p_{1:G}$ of thinned-DDPs obtained by thinning a common Dirichlet process $p$ (see Definition 1), here referred to as the “parent process”. Focusing without loss of generality on the first component $p_{1}$, we study its correlation with the parent process $p$. The results below are direct specializations of those in Sections 2 and 3, obtained by replacing $p_{2}$ with $p$, which corresponds to setting $\ell_{2}$ as an infinite sequence of ones.
@@ Appendix F Additional details on the simulation studies
Estimation of the thinned-DDP was performed using the Gibbs sampler outlined in Section 4; for the complete-pooling and no-pooling models, we used a standard blocked Gibbs sampler for DP mixtures. The truncation parameter was set to 100 for all models. Posterior inference for the CAM and the GM-DDP was performed using the R packages SANple (D’Angelo and Denti 2023) and BNPmix (Corradin et al. 2021), respectively. We ran all algorithms for 3,000 iterations, discarding the first 2,000 as burn-in. Figure F1 shows the distribution of the computation time (in logarithmic scale) for the thinned-DDP, the CAM, and the GM-DDP.
@@ F.1 Additional figures
Figure F3 displays a summary of the distribution of the average length of the pointwise 95% HPD credible intervals for the thinned-DDP, the complete pooling, and no pooling mixtures. Figure F3 displays the same quantity for the thinned-DDP, the GM-DDP, and the CAM.